# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v18

@@ 1 Introduction
  • Calculating the reliability $\alpha\in[0,1]$ of $f\pm\delta f$ seems completely new.
@@
Conventional floating-point arithmetic [6][7][8] computes only the result value $f$. Because $f$ contains unknown amount of rounding error [9][10][11], a floating-point representation with $10^{-16}$ resolution may not be good enough for inputs with $10^{-2}$ to $10^{-6}$ precision. However, statistical Taylor expansion can account for rounding errors as part of the result deviation $\delta f$.
@@
  • Section 7 discusses the evaluation of variance arithmetic on common mathematical library functions, showing the effect of distribution pole.
@@ 2.3 Statistical Taylor Expansion
$\overline{f(x)}$ $=\int_{-\varrho}^{+\kappa}f(x+\tilde{z})\rho(\tilde{z})d\tilde{z}=f(x)+\sum_{n=1}^{\infty}(\delta x)^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n,\kappa);$ (2.5)
$\delta^{2}f(x)$ $=\overline{(f(x)-\overline{f(x)})^{2}}=\overline{f(x)^{2}}-\overline{f(x)}^{2}$
$=\sum_{n=1}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big(\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)\big);$ (2.6)
@@
$\overline{x\pm y}$ $=\zeta(0,\kappa_{x})x\pm\zeta(0,\kappa_{x})y;$ (2.10)
$\delta^{2}(x\pm y)$ $=\zeta(2,\kappa_{x})(\delta x)^{2}+\zeta(2,\kappa_{y})(\delta y)^{2};$ (2.11)
$\overline{xy}$ $=\zeta(0,\kappa_{x})x\;\zeta(0,\kappa_{y})y;$ (2.12)
$\delta^{2}(xy)$ $=\zeta(2,\kappa_{x})(\delta x)^{2}y^{2}+x^{2}\zeta(2,\kappa_{y})(\delta y)^{2}+\zeta(2,\kappa_{x})(\delta x)^{2}\;\zeta(2,\kappa_{y})(\delta y)^{2};$ (2.13)
@@ 2.4 One-Dimensional Examples
$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=1}^{\infty}(\delta x)^{n}\zeta(n,\kappa)\frac{1}{n!};$ (2.14)
@@
$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{\eta(j,x)\eta(n-j,x)}{j!(n-j)!}\big(\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)\big);$ (2.20)
@@ 2.5 Convergence of Variance
Figure 3: When $\kappa=5$, the value of bound moment $\zeta(2n,\kappa)$ (left y-axis) for Gaussian distribution with order $2n$ (x-axis) versus $2N(\kappa)\frac{\kappa^{2n+1}}{2n+1}$ (left y-axis), and their ratio (right y-axis).
@@
$\lim_{n\rightarrow+\infty}$ $\zeta(2n,\kappa)=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.24)
@@ 2.6 Statistical Bounding
Figure 6: Ideal leakage (y-axis) for varying sample count $\kappa_{s}$ (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when $\hat{\kappa}=5$ for Gaussian input uncertainties.
Figure 7: Ideal leakage (y-axis) for varying bound range $\kappa$ (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when $\hat{\kappa}=6$ for Gaussian input uncertainties.
@@
$\epsilon(\kappa)$ $=\omega(\kappa)+2N(\kappa)\kappa;$ (2.30)
@@
Define bounding leakage $\epsilon(\kappa)\equiv 1-\zeta(0,\kappa)$. When the underlying distribution is Normal, Formula (2.30) presents the difference between the linear leakage $\omega(\kappa)$ and the bounding leakage $\epsilon(\kappa)$, while Formulas (2.31) and (2.32) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4]. Figure 4 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$ along the y-axis direction, and that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction. Figure 5 sections Figure 4 along the y-axis direction for $\kappa_{s}=4,5,6$, and shows that to reach the stable variance for a $\kappa_{s}$, the required sample count is $N\gtrsim 10^{3}$, whereas the stable result bounding leakage is smaller for larger bounding range $\kappa$. It also shows that linear leakage converges to 0 with increasing sample count $N$ faster when $\hat{\kappa}$ is larger.
@@
Define ideal leakage as $\varepsilon\equiv 1-\delta^{2}f/\widehat{\delta^{2}}f\simeq\omega$, which quantifies the reliability of $\hat{\delta^{2}}f$ from the underlying distributions and the corresponding sample counts of all inputs, with Formula (2.33) as an example. Figure 6 shows that the stable ideal leakages for the selected functions are less than $10^{-4}$ for $\hat{\delta^{2}}f$, which is small enough for most applications. Figure 7 compares the ideal leakages for $\hat{\delta^{2}}f$ for the selected functions when different $\hat{\kappa}$ is used.
@@ 2.7 Dependency Tracing
When all inputs satisfy the uncorrelated uncertainty assumption, statistical Taylor expansion traces dependencies through the intermediate steps. For example:
@@
  • The dependency tracing of $\delta^{2}f(g(x))$ is demonstrated by $\delta^{2}(f^{-1}(f(x)))=(\delta x)^{2}$. For a reversible transformation such as matrix reversion or FFT (Fast Fourier Transformation), after a round-trip transformation which is a forward transformation followed by a reverse transformation, the original input should be restored.
@@ 2.8 Traditional Execution and Dependency Problem
  • Similarly, conditional executions are often employed to optimize performance and minimize rounding errors, for example, using Gaussian elimination to minimize floating-point rounding errors in matrix inversion [20]. For dependency tracing, such conditional executions should instead be replaced by direct matrix inversion as described in Section 6.
  • Similarly, large calculations are often divided into sequential steps, such as computing $f(g(x))$ as $f(y)|_{y=g(x)}$. This approach also introduces the dependency problem by ignoring dependency tracing within $g(x)$ affecting $f(g(x))$, such as $\overline{(\sqrt{x})^{2}}>\overline{\sqrt{x^{2}}}$ and $\delta^{2}(\sqrt{x})^{2}>\delta^{2}\sqrt{x^{2}}$.
@@ 3.1 Monotonic
Figure 8: Measured upper bound $\delta x$ (left y-axis) for $(1\pm\delta x)^{c}$ across different values of $c$ (x-axis) for Gaussian uncertainty. The corresponding uncertainty bias and uncertainty are also shown (right y-axis). When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure.
Figure 9: Measured upper bound $\delta x$ (left y-axis) for $(1\pm\delta x)^{c}$ across different values of $c$ (x-axis) for uniform uncertainty. The corresponding uncertainty bias and uncertainty are also shown (right y-axis). When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure. When $c\in[0,1]$, uncertainty bias is negative so it does not show up in the plot.
@@ 3.2 Positive
Figure 10: Measured upper bound $\delta x$ (left y-axis) for $\sin(x\pm\delta x)$ across different values of $x$ (x-axis) for Gaussian uncertainty. The corresponding uncertainty bias and uncertainty are also shown (right y-axis). The x-axis is expressed in units of $\pi$.
Figure 11: Measured upper bound $\delta x$ (left y-axis) for $\sin(x\pm\delta x)$ across different values of $x$ (x-axis) for uniform uncertainty. The corresponding uncertainty bias and uncertainty are also shown (right y-axis). The x-axis is expressed in units of $\pi$.
In addition to convergence, the variance expansion may yield negative results, as in Formula (2.20) for $sin(x\pm\delta x)$. Figure 10 shows that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$ for Gaussian input uncertainty. Beyond this upper bound, the result deviation is no longer positive at some expansion order. Conceptually, the upper-bound for $\delta x$ should not be much more than $\pi/2$ to make the result meaningful, which is satisfied as shown in Figure 10. Similar trend holds when the input uncertainty is uniform but with larger upper bound $\delta x$, as shown in Figure 11.
@@ 3.3 Floating-Point Rounding Errors
Variance arithmetic incorporates floating-point rounding errors as the $\delta x$ when converting a floating-point value $x$ into $x\pm\delta x$. Unless all the least 20 bits of the significand of $x$ are zero, $\delta x$ is assumed to be $1/\sqrt{3}$ times of the ULP (Unit in the Last Place [8]) of $x$, because rounding errors are shown to be uniformly distributed within ULP [17].
@@ 4 Verification of Variance Arithmetic
When an input contains unspecified errors, such as numerical errors in library functions or floating-point rounding errors, Gaussian noise with progressively increasing deviations can be added, until ideal coverage is attained. The minimal noise deviation required provides a good estimate of the magnitude of the unspecified input uncertainty deviations. Achieving ideal coverage serves as a necessary verification step to ensure that statistical Taylor expansion has been applied correctly within the given context. The input noise range that yields ideal coverage defines the ideal application range for the analytic function.
@@ 5.3 Distributional Pole
In each case, the transition from $x=0$ to $x=\pm 0.2$ is continuous for the resulting mean, deviation and error distribution.
@@ 6 Matrix Calculations
Figure 15: Error deviations (z-axis) as a function of matrix size (x-axis) and input noise precision (y-axis) for the difference of the two sides of Formula (6.5).
@@
Let $j_{1},j_{2}\dots j_{m}]_{n}$ denote a permutation of length $m$ from the vector $(1,2\dots n)$, and let $j_{1},j_{2}\dots j_{m}>_{n}$ for the ordered permutation of length $m$ [20]. Let $\mathbf{M}$ be a square matrix of size $n$ with element $x_{i,j}\pm\delta x_{i,j},i,j=1,2,\dots n$ at row index $i$ and column index $j$. Fourmula (6.2) presents the Taylor expansion of the determinant for matrix $\mathbf{M}$ when the uncertainties of matrix elements are all independent of each other, whereas Formulas (6.3) and (6.4) gives the statistical Taylor expansion of the determinant $|\mathbf{M}|$ [20]. In Formula (6.4), $\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}$ is a sub-matrix for $\mathbf{M}$, in which $i_{1}\dots i_{m}>_{n}$ contains the row indices, and $j_{1}\dots j_{m}]_{n}$ contains the column indices [20].
[…]

### v19

@@ 1 Introduction
  • Calculating the reliability $[0,1]$ of $f\pm\delta f$ from input uncertainty distributions and sample counts seems completely new.
@@
Conventional floating-point arithmetic [6][7][8] computes only the result value $f$. Because $f$ contains unknown amount of rounding error [9][10][11], a 32-bit floating-point representation with $10^{-7}$ resolution may not be good enough for inputs with $10^{-2}$ to $10^{-6}$ precision. However, statistical Taylor expansion can account for rounding errors as part of the result deviation $\delta f$.
@@
  • Section 7 discusses the evaluation of variance arithmetic on common mathematical library functions, showing the effect of distributional pole.
@@ 2.3 Statistical Taylor Expansion
$\overline{f(x)}$ $=f(x)+\sum_{n=1}^{\infty}(\delta x)^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n,\kappa);$ (2.5)
$\delta^{2}f(x)$ $=\sum_{n=1}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big(\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)\big);$ (2.6)
@@
$\overline{x\pm y}$ $=\zeta_{x}(0,\kappa_{x})x\pm\zeta_{y}(0,\kappa_{x})y;$ (2.10)
$\delta^{2}(x\pm y)$ $=\zeta_{x}(2,\kappa_{x})(\delta x)^{2}+\zeta_{y}(2,\kappa_{y})(\delta y)^{2};$ (2.11)
$\overline{xy}$ $=\zeta_{x}(0,\kappa_{x})x\;\zeta_{y}(0,\kappa_{y})y;$ (2.12)
$\delta^{2}(xy)$ $=\zeta_{x}(2,\kappa_{x})(\delta x)^{2}y^{2}+x^{2}\zeta_{y}(2,\kappa_{y})(\delta y)^{2}+\zeta_{x}(2,\kappa_{x})(\delta x)^{2}\;\zeta_{y}(2,\kappa_{y})(\delta y)^{2};$ (2.13)
@@ 2.4 One-Dimensional Examples
$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=1}^{\infty}(\delta x)^{n}\frac{\zeta(n,\kappa)}{n!};$ (2.14)
@@
$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\eta(j,x)\eta(n-j,x)\frac{\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)}{j!(n-j)!};$ (2.20)
@@ 2.5 Convergence of Variance
Figure 3: When $\kappa=5$, the value of bound moment $\zeta(2n,\kappa)$ (left y-axis) for Gaussian distribution with order $2n$ (x-axis) versus $2N(\kappa)\frac{\kappa^{2n+1}}{2n+1}$ (left y-axis), and their ratio (right y-axis). The value of bound moment $\zeta(2n,\kappa)$ for Uniform distribution is also provided as an comparison.
@@
[…]
@@ 2.6 Statistical Bounding
Figure 6: Ideal leakage $\varepsilon$ (y-axis) for varying sample count $\kappa_{s}$ (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when $\hat{\kappa}=5$ for Gaussian input uncertainties. When $f(x)=x$, ideal leakage equals linear leakage $\omega$.
Figure 7: Ideal leakage $\varepsilon$ (y-axis) for varying sample count $\kappa_{s}$ (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) for Uniform input uncertainties. When $f(x)=x$, ideal leakage equals linear leakage $\omega$.
@@
$\epsilon(\kappa)$ $=\omega(\kappa)+2\rho(\kappa)\kappa;$ (2.30)
@@
Define bounding leakage $\epsilon(\kappa)\equiv 1-\zeta(0,\kappa)$. When the underlying distribution is Normal, Formula (2.30) presents the difference between the linear leakage $\omega(\kappa)$ and the bounding leakage $\epsilon(\kappa)$, whereas Formulas (2.31) and (2.32) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4]. Figure 4 shows that $\omega(\kappa_{s})<\omega(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\omega(\kappa_{s},N)=\omega(\kappa_{s})$ along the y-axis direction, and that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction. Figure 5 sections Figure 4 along the y-axis direction for $\kappa_{s}=4,5,6$, and shows that to reach the stable variance for a $\kappa_{s}$, the required sample count is $N\gtrsim 10^{3}$, whereas the stable resulting linear leakage $\omega$ is smaller for larger measuring bounding range $\kappa_{s}$. It also shows that linear leakage converges to 0 with increasing sample count $N$ faster when $\kappa_{s}$ is larger.
@@
Define ideal leakage as $\varepsilon\equiv 1-\delta^{2}f/\widehat{\delta^{2}}f\simeq\omega$, which quantifies the reliability of $\hat{\delta^{2}}f$ from the underlying distributions and the corresponding sample counts of all inputs, with Formula (2.33) as an example. Figure 6 compares ideal leakage $\varepsilon$ and linear leakage $\omega$ for the selected functions when the input uncertainty distribution is Gaussian, and shows that the example stable ideal leakages are less than $10^{-4}$ when $\hat{\kappa}=5$, which is small enough for most applications.
The unbiased bounding range $\kappa_{min},\kappa_{max}]$ of Uniform distribution is estimated as $1+\frac{1}{N}$ times of the minimal and maximal values in sampled data set. The bonding leakage $\epsilon$ is the percentage of $\kappa_{min},\kappa_{max}]$ outside the true bounding range $\mu-\sqrt{3}\sigma,\mu+\sqrt{3}\sigma]$. Because $\zeta(2n)$ for Uniform distribution increases much slower than $\zeta(2n,5)$ for Gaussian distribution, Figure 7 shows that for the selected functions, ideal leakage equals linear leakage for Uniform input uncertainty, and both decrease linearly with sample count $N$. The difference between Figures 6 and 7 shows the effect of input uncertainty distribution.
@@ 2.7 Dependency Tracing
When all inputs satisfy the uncorrelated uncertainty condition, statistical Taylor expansion traces dependencies through the intermediate steps. For example:
@@
  • The dependency tracing of $\delta^{2}f(g(x))$ is demonstrated by $\delta^{2}(f^{-1}(f(x)))=(\delta x)^{2}$. For a reversible transformation such as matrix reversion or FFT (Fast Fourier Transformation), after a round-trip transformation which is a forward transformation followed by a reverse transformation, the original inputs should be restored.
@@ 2.8 Traditional Execution and Dependency Problem
  • Large calculations are often divided into sequential steps, such as computing $f(g(x))$ as $f(y)|_{y=g(x)}$. This approach also introduces the dependency problem by ignoring dependency tracing within $g(x)$ affecting $f(g(x))$, such as $\overline{(\sqrt{x})^{2}}>\overline{\sqrt{x^{2}}}>\overline{x}$ and $\delta^{2}(\sqrt{x})^{2}>\delta^{2}\sqrt{x^{2}}>\delta^{2}x$.
  • Conditional executions are often employed to optimize performance and minimize rounding errors, for example, using Gaussian elimination to minimize floating-point rounding errors in matrix inversion [20]. For dependency tracing, such conditional executions should instead be replaced by direct matrix inversion as described in Section 6.
@@ 3.1 Monotonic
Figure 8: Measured upper bound $\delta x$ (left y-axis) for $(1\pm\delta x)^{c}$ across different values of $c$ (x-axis) for Gaussian uncertainty. The corresponding resulting uncertainty bias and deviation are also shown (right y-axis). When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure.
Figure 9: Measured upper bound $\delta x$ (left y-axis) for $(1\pm\delta x)^{c}$ across different values of $c$ (x-axis) for uniform uncertainty. The corresponding resulting uncertainty bias and deviation are also shown (right y-axis). When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure. When $c\in[0,1]$, uncertainty bias is negative so it does not show up in the plot.
@@ 3.2 Positive
Figure 10: Measured upper bound $\delta x$ (left y-axis) for $\sin(x\pm\delta x)$ across different values of $x$ (x-axis) for Gaussian uncertainty. The corresponding resulting uncertainty bias and deviation are also shown (right y-axis). The x-axis is expressed in units of $\pi$.
Figure 11: Measured upper bound $\delta x$ (left y-axis) for $\sin(x\pm\delta x)$ across different values of $x$ (x-axis) for uniform uncertainty. The corresponding resulting uncertainty bias and deviation are also shown (right y-axis). The x-axis is expressed in units of $\pi$.
In addition to convergence, the variance expansion may yield negative results when the expansion is terminated arbitrarily, as in Formula (2.20) for $sin(x\pm\delta x)$. Figure 10 shows that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$ for Gaussian input uncertainty. Beyond this upper bound, the result deviation is no longer positive at some expansion order. Conceptually, the upper-bound for $\delta x$ should not be much more than $\pi/2$ to make the result meaningful, which is satisfied as shown in Figure 10. Similar trend holds when the input uncertainty is uniform but with larger upper bound $\delta x$, as shown in Figure 11.
@@ 3.3 Floating-Point Rounding Errors
Variance arithmetic incorporates floating-point rounding errors as $\delta x$ when converting a floating-point value $x$ into $x\pm\delta x$. Unless all the least 20 bits of the significand of $x$ are zero, $\delta x$ is assumed to be $1/\sqrt{3}$ times of the ULP (Unit in the Last Place [8]) of $x$, because rounding errors are shown to be uniformly distributed within ULP [17].
@@ 4 Verification of Variance Arithmetic
When an input contains unspecified errors, such as numerical errors in library functions or floating-point rounding errors, Gaussian noise with progressively increasing deviations can be added to inputs, until ideal coverage is attained. The minimal noise deviation required provides a good estimate of the magnitude of the unspecified input uncertainty deviations. Achieving ideal coverage serves as a necessary verification step to ensure that statistical Taylor expansion has been applied correctly within the given context. The input noise range that yields ideal coverage defines the ideal application range for the analytic function.
@@ 6 Matrix Calculations
Figure 15: Error deviations (z-axis) as a function of input noise precision (x-axis) and matrix size (y-axis) for the difference of the two sides of Formula (6.5).
@@
Let $j_{1},j_{2}\dots j_{m}]_{n}$ denote a permutation of length $m$ from the vector $(1,2\dots n)$, and let $j_{1},j_{2}\dots j_{m}>_{n}$ for the ordered permutation of length $m$ [20]. Let $\mathbf{M}$ be a square matrix of size $n$ with element $x_{i,j}\pm\delta x_{i,j},i,j=1,2,\dots n$ at row index $i$ and column index $j$. Fourmulas (6.1) and (6.2) present the Taylor expansion of the determinant for matrix $\mathbf{M}$ when the uncertainties of matrix elements are all independent of each other, whereas Formulas (6.3) and (6.4) gives the statistical Taylor expansion of the determinant $|\mathbf{M}|$. In Formula (6.4), $\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}$ is a sub-matrix for $\mathbf{M}$, in which $i_{1}\dots i_{m}>_{n}$ contains the row indices, and $j_{1}\dots j_{m}]_{n}$ contains the column indices [20].
The square matrix whose element is $(-1)^{i+j}|M_{j,i}|$ is defined as the adjugate matrix [20] $\mathbf{M}^{A}$ to the original square matrix $\mathbf{M}$. By assigning a random integer between $2^{8},+2^{8}]$ to each element, $\mathbf{M}$ is created such that $\mathbf{M}^{A}$ can be calculated precisely using floating-point arithmetic when $n\leq 7$. By adding Gaussian noise $\tilde{x}_{i,j}$ of deviation $\delta x$ to each element of $\mathbf{M}$, $\widetilde{\mathbf{M}}$ is created as shown in Formula (6.1). The value error of $\mathbf{M}^{A}$ is the difference between $\widetilde{\mathbf{M}^{A}}$ and $\mathbf{M}^{A}$, whereas the result deviation is calculated using Formula (6.4), such that $\widetilde{\mathbf{M}^{A}}-\mathbf{M}^{A}$ is a distribution test. Figure 14 shows that the error deviations of $\widetilde{\mathbf{M}^{A}}-\mathbf{M}^{A}$ are very close to $1$ except when the matrix size is $8$ and the input uncertainty is $\delta x=10^{-17}$. This abnormality is caused by floating-point rounding errors only at this point, to result in a error deviation of $3.2$. Figure 14 demonstrates a typical distribution test of error deviation versus input noise deviation $\delta x$ and another algorithm specific dimension, with ideal coverage for all region except when $\delta x\rightarrow 0$.